A 66kW broadcasting antenna emits radio waves at a frequency of 1 MHz. How many photons are emitted per second?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
number of photons emitted per second = 
 
n = Power / Energy of Photons
 
 = P/ hv
 
 = 66000/ (6.634* 10^-34 * 10^6) = 9948 * 10^28 = 10^31photons approx